Grade upto college level Physical Chemistry

A sample of hydrazine suphate (N2H6SO4) was dissolved in 100 ml. of water, 10 ml of this solution was reacted with excess of ferric chloride solution and warmed to complete the reaction. Ferrous ion formed was estimated and it required 20 ml. of M/50 potassium permanganate solution. Estimate the amount of hydrazine sulphate in one litre of the solution.
Reaction :
4Fe3+ + N2H4 → N2 + 4Fe2+ + 4H+
MnO-4 + 5 Fe2+ + 8H+ → Mn2+ + 5Fe3+ + 4H2O

Amit Saxena
12 Years agoGrade upto college level

1 Answer

Navjyot Kalra
12 Years ago
No. of equivalent of KMnO4
= No. of equivalents of hydrazine sulphate.
N2H4 → N2
2 0
Change in oxidation state for each N2H4 = 2 * 2 = 4
Equivalent weight of N2H6SO4 = 130/4 = 32.5
Normality of KMnO4 5 * 450 (∵ valence factor = 5)
Number of equivalents of KMnO4 = 20 * 5/50 *1000 = 1/500 and if weight of hydrazine sulphate be x gm then equivalents of hydrazine sulphate = x/32.5
∴ 1/500 = x/32.5 or x 32.5/500 = 0.065 g
Hence wt. of N2H6SO4 in 10 ml solution = 0.065 g
∴ Wt. of N2H6SO4 in 1000 ml solution = 6.5 g